2. If • a<b<c • Sºf(x) dx = 5 • Sog(x) = 3 • Sg(x) dx = 7 Then, what is the value of f(x) - g(x)]dx?
Answer #1

s poders Z ♡ = 7-32. 4 [Troy - swa) dut el post - proste = 5-4 = 12